OpenStudy (anonymous):

A rectangle has a perimeter of 26 feet. Twice its length (l) is one foot more than three times its width (w). What is the length of the rectangle, in feet?

OpenStudy (anonymous):

perimeter 2(l+w)=26 l+w=13 ....(1) again 2l=w+1 w=2 l-1 put in (1) and find the value of l and then w.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw